# A Particle Algorithm for Mean-Field Variational Inference

### v2

@@ 4 Theoretical analysis
Let $h=1/(LN^{1/4})$ and $B=\lceil 1/(Lh)\rceil$. We have
$\sqrt{\mathbb{E}\mathcal{W}_{2}^{2}(q_{\bm{X}_{n}},q_{*})}$ $\leq(1-\alpha h/4)^{n}\sqrt{\mathbb{E}\mathcal{W}_{2}^{2}(q_{\bm{X}_{0}},q_{*})}+\frac{\sqrt{m}}{N^{1/4}}\cdot\frac{CL}{\alpha^{3/2}}\bigg{(}\sqrt{\log N}+\sqrt{1+\widetilde{L}^{2}\alpha/L^{4}}\bigg{)},$
where $C>0$ is a universal constant.
We note that the non-asymptotic upper bound in Corollary 4.1 holds for any number of particles $N$ and any number of iterations $n$, and it converges to 00 as $N$ and $n$ grow at appropriate rates (see below for a detailed analysis of the required sizes of $N$ and $n$ to achieve a given precision $\varepsilon\in(0,1)$). Thus, our result establishes the long-time convergence of Algorithm 2. To the best of our knowledge, this is the first such convergence guarantee for particle-based MFVI. The particle algorithm proposed in Yao and Yang (2022) has cumulative numerical errors that do not diminish even when the number of particles and the number of iterations grow to infinity. Furthermore, the theoretical analysis therein does not address errors arising from the particle approximation.
By Corollary 4.1, if we take the step-size $h=1/(LN^{1/4})$ and the batch size $B=\lceil 1/(Lh)\rceil$, then for any $\varepsilon\in(0,1)$, as long as the number of particles $N$ satisfies $N\geq Km^{2}(\log(m/\varepsilon^{2}))^{2}\varepsilon^{-4}$ (where $K$ is a sufficiently large constant depending only on $\alpha,L,\tilde{L}$), the number of iterations $T$ required in order for $\sqrt{\mathbb{E}\mathcal{W}_{2}^{2}(q_{\bm{X}_{T}},q_{*})}\leq\varepsilon$ is
$\mathcal{O}_{\alpha,L}\bigg{(}N^{1/4}\max\Big{\{}\log\Big{(}\sqrt{\mathbb{E}\mathcal{W}_{2}^{2}(q_{\bm{X}_{0}},q_{*})}\varepsilon^{-1}\Big{)},1\Big{\}}\bigg{)}.$
In particular, if the number of particles $N$ is taken to be of order $m^{2}(\log(m/\varepsilon^{2}))^{2}\varepsilon^{-4}$, the required number of iterations is
$\mathcal{O}_{\alpha,L,\tilde{L}}\bigg{(}\sqrt{m\log(m/\epsilon^{2})}\varepsilon^{-1}\max\Big{\{}\log\Big{(}\sqrt{\mathbb{E}\mathcal{W}_{2}^{2}(q_{\bm{X}_{0}},q_{*})}\varepsilon^{-1}\Big{)},1\Big{\}}\bigg{)}.$ (4.2)
Notably, the requirement (4.2) on the number of iterations grows mildly with the dimension $m$ and does not suffer from the curse of dimensionality.
We also note that our algorithm, Algorithm 2, provides a direct approximation $q_{\bm{X}_{n}}$ to the MFVI solution $q_{*}$, with its approximation accuracy guaranteed by Theorem 4.1 and Corollary 4.1. Langevin Monte Carlo (LMC), by contrast, tracks the evolution of a single particle. Theoretical results for LMC (Dalalyan, 2017b; Durmus and Moulines, 2017; Cheng and Bartlett, 2018; Durmus et al., 2019; Durmus and Moulines, 2019) primarily focus on bounding the $\mathcal{W}_{2}$ distance (or other metrics on probability measures) between the distribution of this particle at the $n$-th iteration and the target distribution. In order to approximate the target distribution using the output of LMC, one needs to generate many such particles and use their empirical distribution. One option is to run $T_{0}$ independent replications of LMC, each time collecting the particle $\mathbf{x}_{t,n}\in\mathbb{R}^{m}$ (where $t\in[T_{0}]$) at the $n$-th iteration, and use the empirical measure $T_{0}^{-1}\sum_{t=1}^{T_{0}}\delta_{\mathbf{x}_{t,n}}$ to approximate the target distribution. Note that this approach would suffer from the curse of dimensionality: in high dimensions, the $\mathcal{W}_{2}$ distance between the empirical distribution and the true one is known to converge very slowly as the sample size $T_{0}$ increases (Weed and Bach, 2019).
5 Proof of Theorem 4.1
In this section, we present the proof of Theorem 4.1.
5.1 Preparations
The main idea is to couple the iterates of Algorithm 2 with a stationary process. First, we introduce a metric for measuring the difference between two particle swarms.
Definition 5.1.
For two Borel probability distributions $\mu$ and $\nu$ on $\mathbb{R}^{m\times N}$, define $\mathcal{C}(\mu,\nu)$ as the set of all couplings between $\mu$ and $\nu$. Let
$\widetilde{\mathcal{W}}(\mu,\nu)=\inf_{\gamma\in\mathcal{C}(\mu,\nu)}\sqrt{\mathbb{E}_{(\bm{X},\bm{Y})\sim\gamma}\mathcal{W}_{2}^{2}(q_{\bm{X}},q_{\bm{Y}})}\,.$
Let $\bm{Z}$ and $\bm{W}$ be two random arrays in $\mathbb{R}^{m\times N}$, which induce random product distributions $q_{\bm{Z}}$ and $q_{\bm{W}}$ as in Definition 3.2, respectively. Intuitively, if $q_{\bm{W}}$ concentrates around a fixed probability distribution $\mu$, then $\mathbb{E}\mathcal{W}_{2}^{2}(q_{\bm{Z}},q_{\bm{W}})$ should be close to $\mathbb{E}\mathcal{W}_{2}^{2}(q_{\bm{Z}},\mu)$ and thus the former distance is insensitive to the dependence between $\bm{Z}$ and $\bm{W}$. In that case, $\sqrt{\mathbb{E}\mathcal{W}_{2}^{2}(q_{\bm{Z}},q_{\bm{W}})}$ is close to the infimum of $\widetilde{\mathcal{W}}\left(\mathrm{Law}(\bm{Z}),\mathrm{Law}(\bm{W})\right)$ over all couplings between $\mathrm{Law}(\bm{Z})$ and $\mathrm{Law}(\bm{W})$. This is confirmed by the lemma below.

### v3

@@ 4 Theoretical analysis
Let Assumption 4.1 hold. Run Algorithm 2 with $N\geq 9$, $h=\frac{\alpha\log N}{L^{2}N}$ and $B=1$. Suppose that the columns of $\bm{X}_{0}\in\mathbb{R}^{m\times N}$ are drawn independently from a product distribution $q_{0}$, whose marginals are log-concave with variance bounded by $\alpha$. There exists a universal constant $C>0$ such that for any $n\in\mathbb{N}$,
$\sqrt{\mathbb{E}\mathcal{W}_{2}^{2}(q_{\bm{X}_{n}},q_{*})}$ $\leq\exp\bigg(-\frac{n\alpha^{2}\log N}{4L^{2}N}\bigg)\mathcal{W}_{2}(q_{0},q_{*})+C\sqrt{\frac{m\log N}{\alpha N}}.$
Suppose that $\alpha$ and $L$ are constants. Choose any $\varepsilon\in(0,1)$. There exists a constant $K$ determined by $\alpha$ and $L$, such that when $N>Km\varepsilon^{-2}\log(m\varepsilon^{-2})$, we have $C\sqrt{\frac{m\log N}{\alpha N}}\leq\varepsilon/2$. Meanwhile, when
$n\geq\frac{4L^{2}N}{\alpha^{2}\log N}\log[2\varepsilon^{-1}\mathcal{W}_{2}(q_{0},q_{*})],$
we have $\exp(-\frac{n\alpha^{2}\log N}{4L^{2}N})\mathcal{W}_{2}(q_{0},q_{*})\leq\varepsilon/2$. Therefore, taking
$N=O_{\alpha,L}\Big(m\varepsilon^{-2}\log(m\varepsilon^{-2})\Big)\quad\text{and}\quad n=O_{\alpha,L}\Big(m\varepsilon^{-2}\log[2\varepsilon^{-1}\mathcal{W}_{2}(q_{0},q_{*})]\Big)$
with sufficiently large constant factors can guarantee $\sqrt{\mathbb{E}\mathcal{W}_{2}^{2}(q_{\bm{X}_{n}},q_{*})}\leq\varepsilon$. Both $N$ and $n$ have linear dependence on $m$ up to logarithmic factors.
We conclude this section with a comparison between theoretical guarantees for Algorithm 2 and Langevin Monte Carlo (LMC).
  • Our algorithm maintains $N$ interacting particles. According to Corollary 4.1, it directly approximates the MFVI solution $q_{*}$ up to $O(\sqrt{\frac{m\log N}{N}})$ in $\mathcal{W}_{2}$. The fast convergence of particle approximation could compensate for the discrepancy between $q_{*}$ and the target distribution $p$.
  • LMC tracks the evolution of a single particle. Theoretical results (dalalyan2017theoretical; durmus2017nonasymptotic; cheng2018convergence; durmus2019analysis) quantify how its distribution converges to $p$. In order to approximate $p$, one generates a large number of such particles (e.g., by running $N$ independent replicates of LMC) and uses their empirical distribution. This would suffer from the curse of dimension: if we only assume Assumption 4.1, then the empirical distribution formed by $N$ independent particles converges in $\mathcal{W}_{2}$ at a rather slow rate of $N^{-1/m}$ when the dimension $m$ exceeds 3 (FGu15).
5 Proofs
We present the proofs of our main results.
5.1 Proof of Theorem˜4.1
We begin with a recursive bound based on a coupling between the iterates $\{\bm{X}_{n}\}_{n=0}^{\infty}$ of Algorithm˜2 and a stationary process. The proof is deferred to Section˜5.5.